OpenStudy (anonymous):

When planning his crops, Farmer Sue knows that her 15 acres can support apples and pecans. She wants to make $8,500 from her crops. She can make $1,050 per acre of apples (the variable a) and $2,500 per acre of pecans (the variable p). Which equation below would be a constraint in her system of equations?

OpenStudy (dmb1215):

i know it is p+a=8,500

OpenStudy (dmb1215):

but do we put the 15 with any of them

OpenStudy (***[isuru]***):

ok... they say that the total Sue expects from her crops is = $8500 and she gets $1050 per acre from apples... if she grows "a" acres then her outcome will be 1050a she gets $2500 per acre from pecans if she grows "p" acres then the outcome will be 2500p but since her total is 8500 1050a + 2500p = 8500 ----(1) 21a + 50p = 170 -----(1) since her land area is 15 acres a + p = 15 ----(2) so.. ur choices will contain one out of the followings.. hope this will help ya :)